### v3

# PerturBench: Benchmarking Machine Learning Models for Cellular Perturbation Analysis

Abstract
We introduce a comprehensive framework for perturbation response modeling in single cells, aimed at standardizing benchmarking in this rapidly evolving field. Our approach includes a modular and user-friendly model development and evaluation platform, a collection of diverse perturbational datasets, and a set of metrics designed to fairly compare models and dissect their performance nuances. Through extensive evaluation of both published and baseline models across diverse datasets, we highlight the limitations of widely used models, such as mode collapse. We also demonstrate the importance of rank metrics which complement traditional model fit measures, such as RMSE, for validating model effectiveness. Notably, our results show that while no single model architecture clearly outperforms others, simpler architectures are generally competitive and scale well with larger datasets. Overall, this benchmarking exercise sets new standards for model evaluation, supports robust model development, and advances the potential of these models to use high-throughput genetic and chemical screens for disease target discovery.
1 Introduction
Perturbing biological systems such as cells using small molecules and genetic modifications can help researchers to uncover causal drivers of diseases and identify potential therapeutic targets [8, 27, 11]. Advances in CRISPR technology and lab automation have enabled these experiments, which we refer to as perturbation screens, to be conducted at scale with up to hundreds of thousands of perturbations applied in parallel in a single experiment [49]. These perturbation screens have been combined with modern RNA-sequencing technology to measure gene expression profiles at single cell resolution, creating an atlas of cellular snapshots that reveal perturbation effects [36, 54, 37, 1, 15, 8, 51, 61].
However, measuring the perturbation effects of all roughly $20,000$ protein coding genes or $10^{60}$ drug-like chemicals remains prohibitively expensive, especially when combinations of perturbations and multiple tissues/cell types are considered [44, 45]. As a result, there has been a growing interest in generative machine learning approaches that can predict the effects of perturbations on gene expression.
Specifically, researchers have developed models that can generate counterfactual, out of sample (oos) predictions of perturbation effects [19]. One use case, which we call covariate transfer, involves training a model on perturbation effects measured in a set of covariates (e.g., cell lines) and predicting those effects in another covariate where they have not been observed. Another use case — combo prediction, involves training a model on individual perturbation effects and predicting the effects of multiple perturbations in combination. The ultimate goal of these models is to enable in-silico screening across the vast space of unobserved perturbations for novel therapeutic discovery.
Comparing the performance of published models has unfortunately remained a challenge due to inconsistent benchmarking datasets and metrics. The NeurIPS 2023 perturbation prediction challenge provided a novel chemical perturbation, but was limited to just that dataset with the covariate transfer task and a single metric [10]. Additionally, the challenge did not benchmark published models. The sc-perturb database provides datasets with unified metadata, but does not attempt to benchmark models [41]. Ahlmann-Eltze et al., [2], Wenteler et al., [56], Csendes et al., [12], and Wong et al., [58] assessed only the performance of single cell foundation models and GEARS and all studies used only MSE as an evaluation metric, which may not capture key aspects of model performance. Kernfeld et al., [28] focused on using gene regulatory networks to predict perturbation response prediction. Li et al. [31, 32] performed a more comprehensive assessment of published deep learning models across multiple datasets, but were missing strong deep learning baselines and used a limited set of metrics. To our knowledge, none of the existing benchmarking studies offer a model development framework to build perturbation response prediction models. The existing studies also do not perform ablation studies to identify the specific architectural components that drive model performance.
In this work, we provide a rigorous, quantitative assessment of the field of perturbation response modeling by: (1) introducing a highly modular and user-friendly framework in the form of a codebase for model development and evaluation, (2) curating diverse perturbational datasets and defining biologically relevant tasks, (3) defining a set of metrics that enable comparison of different models and capture key failure modes, (4) performing extensive evaluation of published perturbation models, strong baselines, and individual model components. Figure 1 illustrates our approach.
Figure 1: A) Single cell perturbational datasets at multiple scales. B) Biologically relevant covariate transfer and combinatorial prediction data splits. C) Dataloaders support two training strategies: 1) control matching which involves mapping a control cell to a perturbed cell and 2) disentanglement which maps a perturbed cell to itself. D) A model zoo with modular components such as relevant baseline models, adversarial loss, perturbation sparsity, and others. E) Standardized benchmarking suite supporting flexible pipelines and metrics for evaluating models
We reproduce key components of published models that cover a spectrum of architectures, and evaluate them against each other and strong baseline models. We specifically test the models on difficult tasks, simulating how they will be deployed in real-life contexts. Our findings reveal that some widely used models are prone to “mode” or “posterior” collapse (see Appendix 9.3 for more details). Since a common use-case of these models is to run in-silico screens that rank perturbations by a desired effect (e.g. reversing a disease state) [55], we propose rank metrics complementary to traditional measures of model fit (e.g. root mean squared error (RMSE)) that specifically assess the models’ ability to accurately order perturbations and act as a safeguards against model collapse. In addition, we demonstrate that models with simple architectures can outperform some of the more sophisticated models, and scale well with larger datasets.
We anticipate that our codebase, together with this benchmark and the accompanying metrics, will serve as a valuable framework for the community to develop more robust perturbation response prediction models. The PerturBench library can be found on GitHub at https://github.com/altoslabs/perturbench/.
2 Datasets and Tasks
Many published models have been evaluated on relatively small datasets, where most of the data are seen in the train split. However, in real-world settings, we often have complex datasets that only contain a small fraction of the perturbation effects we are interested in predicting. Thus, we select datasets and create tasks that mirror these real-world challenges. We select five published datasets, Norman et al., [39], Srivatsan et al., [51], Frangieh et al., [18], McFaline-Figueroa et al., [38], and Jiang et al., 2024a [24] which include at least 100 perturbations and cover a diversity of perturbation modalities (chemical vs genetic), combinatorial perturbations, dataset sizes, and covariates. We provide a cursory overview in Table 1 and more information in Appendix 10.1. Here, we define a biological state as a unique set of covariates that we plan to model (e.g. cell line). Details of the preprocessing of the datasets are in the Appendix 10.3.
Table 1: Summary of benchmarking datasets.
Dataset Single Dual Modality Biological Cells Task
perturbations perturbations states
Srivatsan20 188 0 chemical 3 178,213 covariate transfer
Frangieh21 248 0 genetic 3 218,331 covariate transfer
Jiang24 219 0 genetic 30 1,628,476 covariate transfer
McFalineFigueroa23 525 0 genetic 15 892,800 covariate transfer
Norman19 155 131 genetic 1 91,168 combo prediction
We create a covariate transfer split for the Srivatsan20, Frangieh21 and Jiang24 datasets as well as a combo prediction task for the Norman19 dataset. In addition, we study two scenarios: data scaling and imbalanced data. In the former, we benchmark model performance with increasing training data. In the latter, we simulate increasing the imbalance of perturbations observed in different covariates. Details of the experiments are in sections 5.3 and 9.2 respectively. The aim of both scenarios is to simulate how models will be deployed in practice, where there are often complex covariates, imbalanced datasets, and/or large amounts of missing data [see e.g. 16]. Additional details about data splitting implementation can be found in the Appendix 10.4.
3 Perturbation Prediction Models
3.1 Modeling counterfactuals
Perturbation response models aim to predict out-of-sample effects of genetic or chemical interventions on cells. Here, we define out of sample as predicting effects in unobserved covariates or unobserved perturbation combinations. However, RNA sequencing technology destroys the cell, making it impossible to observe its gene expression state before and after perturbation. Published models use two main strategies to learn representations of perturbation effects: matching methods to match control and perturbed cells, or disentanglement strategies within autoencoder architectures to separate the effects of perturbations from the baseline cell state.
Matched Controls
Matching treated outcomes to controls is a common approach to identify treatment effects [see e.g. 52, Section 1.3 for a historical summary]. In the context of perturbation effect prediction, matching control and perturbed cells has been used by a variety of published models such as GEARS [47], scGPT [13], and scFoundation [21]. For the matching approach, ensuring that the control cell is from the same cell type, experiment or batch as the perturbed cells helps reduce potential confounding effects, but cannot account for unobserved sources of variance. A more complex approach is to use optimal transport to identify the control cell most likely to transition into a given perturbed cell [25, 9, 29]. This enables prediction of the full distribution of cellular responses, instead of just the average response.
Disentanglement
An alternative to matching methods involves disentanglement [5], which enables models to separate the unperturbed cellular state and the perturbation effect. The compositional perturbation autoencoder (CPA) [35] uses an adversarial classifier to ensure that the unperturbed “basal” state is free of any perturbational information, forcing the perturbation encoder to learn a meaningful representation of the perturbation. These representations are added to control cell encodings during inference to generate counterfactual predictions. Biolord [42] partitions the latent space into subspaces and optimizes those latent spaces to represent covariates and perturbations, which can be recombined during inference to generate counterfactual predictions. sVAE [34] leverages recent results by Lachapelle et al., [30] demonstrating that enforcing a sparsity constraint can induce disentanglement. Bereket and Karaletsos, [6] build on sVAE using an additive conditioning for the perturbations and demonstrate that their model, SAMS-VAE, offers biological interpretability of the latent encodings.
3.2 Models for benchmarking
In this paper, we implement a range of perturbation response models with diverse architectures such as CPA, Biolord, SAMS-VAE and GEARS. Our aim is to assess the core modeling assumption behind these models, such as the adversarial classifier in CPA and the sparse additive mechanism in SAMS-VAE. See Appendix 10.5 for details. These models are marked with following the model name (e.g. CPA).
To investigate the effect of disentanglement, we ablate the adversarial component from CPA and refer to the new version as CPA (noAdv). The rest of the model is unchanged, that is despite the basal state of cells (i.e. $z_{basal}$) are contaminated with perturbation and covariate information, counterfactual gene expressions are still generated by adding the target perturbation embedding to $z_{basal}$.
Working under the same hypothesis, we removed the binary mask from SAMS-VAE, thus discarding the sparsity assumption which drives the disentanglement and offers interpretability. We also discard the distributional assumption on the perturbation embedding, arriving at a simplified version which we refer to as SAMS-VAE (S) that is free of any global latent variables.
We also experiment with scGPT, a single-cell gene expression foundation model [13], to embed gene expression — as inputs to CPA and our latent additive baseline model. Our aim is to understand whether using embeddings from a pretrained foundation model would improve performance.
3.3 Baseline models
We also implement and benchmark the following baselines:
Linear
The linear baseline model uses the control matching approach. Given a perturbed cell, $x\textquoteright$ ’, we sample a random control cell with matched covariates, $x$, and reconstruct $x^{\prime}$ by applying one linear layer given the perturbation and covariates:
$x^{\prime}=x+f_{\mathrm{linear}}(p_{\mathrm{one\_hot}},cov_{\mathrm{one\_hot}}),$ (1)
where $p_{\mathrm{one\_hot}}$ denotes the one-hot encoding of the perturbation and $cov_{\mathrm{one\_hot}}$ denotes one-hot encodings of covariates (e.g. cell type).
Latent Additive
We extend the linear model into a stronger baseline which we call latent additive, by encoding expression values and perturbations into a latent space $\mathsf{Z}\subseteq\mathbb{R}^{d_{z}}$, i.e.
$z_{\mathrm{ctrl}}=f_{\mathrm{ctrl}}(x),\quad\text{and}\quad z_{\mathrm{pert}}=f_{\mathrm{pert}}(p_{\mathrm{one\_hot}}),$
Subsequently, we reconstruct the expression value by decoding the added latent space representation $x\textquoteright=f_{\mathrm{dec}}(z_{\mathrm{ctrl}}+z_{\mathrm{pert}})$. All functions $f_{\mathrm{dec}},f_{\mathrm{ctrl}},f_{\mathrm{pert}}$ are implemented as multilayer perceptrons (MLPs) with dropout [50] and layer normalization [4].
Decoder Only
We introduce another class of baseline which we refer to as decoder only, that does not leverage gene expression and aims to predict the perturbation effect solely from covariates, or perturbation, or a mix of both. Consequently, model prediction can be expressed as $x^{\prime}=f_{\mathrm{dec}}(z)$ for $z\in\{p_{\mathrm{one\_hot}}\}\cup\{{cov}_{\mathrm{one\_hot}}\}\cup\{(p_{\mathrm{one\_hot}},{cov}_{\mathrm{one\_hot}})\}$. This baseline can be used to establish a performance lower-bound when it does not receive any perturbation information, simulating mode or posterior collapse where the model predicts the same effect for every perturbation.
4 Benchmarking
Existing perturbation response modeling studies use different metrics, making direct comparison between models difficult. Here, we develop a standardized, modular benchmarking suite with a variety of metrics that mimic key downstream applications and capture common model failure modes. See Appendix 7 for implementation details.
4.1 Population Aggregation
Since we only have conditionally paired populations of control cells and perturbed cells, a model has to take unperturbed (i.e. control) cell states and predicts what their states would look like if they had been perturbed. To evaluate these counterfactual predictions, we thus compute aggregate measures from the control and perturbed cell populations, and compare the predicted aggregates to the ground truth aggregates. We use the population mean and log fold-change (LogFC) in this paper, but also implement log p-value and distributional metrics such as MMD with RBF kernel and energy distance in our codebase.
4.2 Metrics
Figure 2: Visualization of the ranking approach. We measure which perturbation in the data is closest to the predicted perturbation as measured by the closeness of their transcriptomes.In case A the rank metric for prediction X is $\mathrm{rank}(X)=\frac{0}{6}=0$, in case B $\mathrm{rank}(Y)=\frac{4}{6}=0.67$.
We select a range of metrics that evaluate different aspects of the predicted perturbation response accuracy. We use RMSE (as recommended in Ji et al., 23) to compare the average predicted cell states to the observed cell states. We also use cosine similarity to assess the similarity of the predicted versus observed perturbation effects (LogFCs). However, as we show in Appendix 9.3, these “global” fit-based metrics fail to fully capture all aspects of a model’s performance. Hence, we introduce a set of rank-based metrics that can be seen as a measure of specificity of the model to different perturbations.
4.2.1 Rank Metrics
Since the space of possible genetic or chemical perturbations is massive, a common application of these models is to rank perturbations by their ability to induce some desired cell state, and select only the top perturbations for experimental validation [55].
Unfortunately, there are no metrics developed to capture this application so far, which results in models that generate predictions with high cosine similarity or low RMSE to the observed gene expression, but fail to capture smaller but key changes that uniquely distinguish the effects of one perturbation from others. In particular, mode or posterior collapse where a model always generates the same prediction irrespective of target perturbation may still result in decent cosine similarity or RMSE on particular dataset.
We find it difficult to use existing information retrieval metrics such as the mean reciprocal rank [43] as that would require a specific desired cell state to create rankings, potentially biasing our evaluation. Therefore, we introduce a rank-based metric that measures, for a given observed perturbation, how close the model prediction is to the observation, compared to predictions made for other perturbations. The rank metric is computed on a per perturbation basis:
$\mathrm{rank}_{\mathrm{average}}$ $:=\frac{1}{p}\sum_{i=1}^{p}\mathrm{rank}(\hat{x}_{i}),\quad\mathrm{rank}(\hat{x}_{i}):=\frac{1}{p-1}\sum_{1\leq j\leq p\atop j\neq i}\mathbb{I}(\mathrm{dist}(\hat{x}_{j},x_{i})\leq\mathrm{dist}(\hat{x}_{i},x_{i})),$ (2)
where $p$ is the number of perturbations that are being modelled, $\hat{x}_{i},x_{i}$ are the predicted and observed (average) expression value of perturbation $i$, and $\mathrm{dist}$ is a generic distance. Figure 2 shows two examples of perturbations predictions and the respective computations of rank metrics. The rank metric is always a number between 00 and $1$, where 00 is a perfect score and $0.5$ is the expected score of a random prediction. Each fit-based metric has a corresponding rank metric.
4.3 Benchmarking Rules
Since model performance often varies with hyperparameters, we run hyperparameter optimization (HPO) for every model in each dataset, task and scenario to ensure accurate model comparisons. Specifically, we use optuna [3] with the default tree-structured Parzen estimator [7]. Each HPO run includes at least 60 trials with 6 trials running in parallel, and we select the best hyperparameters using a combination of the RMSE and RMSE rank metric, i.e. $\mathrm{RMSE}+0.1\cdot\mathrm{rank}_{\mathrm{RMSE}}$. We find this approach results in good overall performance. Additional details are in Appendix 10.6, along with the best hyperparameters for each model/dataset/task. For the best hyperparameter configuration, we run model training four additional times with different seeds to assess stability.
5 Experiments
In this section, we summarize the results of the covariate transfer and combo prediction tasks using different datasets, as well as the data scaling scenario. Additional results, including results for the Frangieh21 dataset, imbalanced data scenario, and implementation details can be found in Appendices 9 and 10.
Table 2: Results of the covariate transfer experiment measuring generalization across cell types in the Srivatsan20 dataset. Mean $\pm$ one standard deviation reported across 5 seeds. Best performance per metric is indicated in bold. The sme convention applies to subsequent tables.
Model Cosine RMSE Cosine RMSE
log fold change (LogFC) mean LogFC rank mean rank
CPA $0.38\pm 6\times 10^{-3}$ $0.020\pm 3\times 10^{-4}$ $0.15\pm 1\times 10^{-2}$ $0.16±\pm 7\times 10^{-3}$
CPA (noAdv) $0.40\pm 5\times 10^{-3}$ $0.020\pm 1\times 10^{-4}$ $\mathbf{0.09\pm 4\times 10^{-3}}$ $\mathbf{0.10\pm 5\times 10^{-3}}$
CPA (scGPT) $0.39\pm 9\times 10^{-3}$ $0.020\pm 3\times 10^{-4}$ $0.13\pm 2\times 10^{-2}$ $0.16\pm 2\times 10^{-2}$
SAMS-VAE $0.44\pm 1\times 10^{-3}$ $0.023\pm 8\times 10^{-5}$ $0.17\pm 1\times 10^{-2}$ $0.17\pm 1\times 10^{-2}$
SAMS-VAE (S) $\mathbf{0.53\pm 1\times 10^{-2}}$ $\mathbf{0.018\pm 3\times 10^{-4}}$ $0.12\pm 2\times 10^{-2}$ $0.13\pm 1\times 10^{-2}$
Biolord $0.18\pm 1\times 10^{-1}$ $0.086\pm 4\times 10^{-2}$ $0.37\pm 2\times 10^{-2}$ $0.35\pm 1\times 10^{-1}$
LA $0.45\pm 2\times 10^{-3}$ $\mathbf{0.018\pm 6\times 10^{-5}}$ $0.13\pm 4\times 10^{-3}$ $0.15\pm 3\times 10^{-3}$
LA (scGPT) $0.50\pm 4\times 10^{-3}$ $\mathbf{0.017\pm 1\times 10^{-4}}$ $0.13\pm 7\times 10^{-3}$ $0.14\pm 5\times 10^{-3}$
Decoder $0.35\pm 5\times 10^{-3}$ $\mathbf{0.018\pm 1\times 10^{-4}}$ $0.16\pm 1\times 10^{-2}$ $0.14\pm 7\times 10^{-3}$
Decoder (Cov) $0.30\pm 1\times 10^{-2}$ $0.023\pm 3\times 10^{-5}$ $0.47\pm 9\times 10^{-3}$ $0.50\pm 4\times 10^{-2}$
Linear $0.16\pm 1\times 10^{-2}$ $0.030\pm 5\times 10^{-4}$ $0.28\pm 5\times 10^{-3}$ $0.27\pm 2\times 10^{-3}$
5.1 Predicting Perturbation Effects Across Cell Types
We begin with the covariate transfer task and assess a model’s ability to predict the effects of drug treatment in cell types where the drugs have not been observed. For each cell line in the Srivatsan20 dataset, we held out 30% of the perturbations for validation and testing, ensuring that any held-out perturbations have been observed in the two other cell types.
The results are summarized in Table 2. Of the three published models, CPA performs best on the rank metrics, while SAMS-VAE is slightly worse. However, SAMS-VAE performs better on the fit based RMSE and cosine LogFC metrics. BioLord underperforms both models on all metrics.
We performed ablation studies by removing the adversarial component from CPA, resulting in the CPA (noAdv). Strikingly, CPA (noAdv) outperforms the original version and is the best performing model overall on the rank metrics. We also ablated the sparse mask and global latent variables from SAMS-VAE, resulting in SAMS-VAE (S), and observe that SAMS-VAE (S) also beats the original implementation on all metrics. These results highlight the need for a modular model development platform that enables ablation studies.
Our decoder-only baseline which does not leverage perturbation information, Decoder (Cov), also performs well on cosine LogFC and RMSE fit, yet does no better than random on rank metrics. This suggests that Decoder (Cov) can find a single expression vector that achieves a decent fit to all perturbations in a given cell type, highlighting the need for our rank metrics that assess whether models can correctly order perturbations. Appendix 9.3 contains a more detailed assessment, where we are able to establish that SAMS-VAE (S) as well as our baselines, are less prone to mode collapse, based on analysis in the Srivatsan20 and Frangieh21 datasets.
By replacing gene expression inputs with scGPT cell embeddings, we are able to see marginal improvement in the performance for CPA and LA.
We also benchmark model performance with the Frangieh21 dataset, where we assess models’ ability to transfer perturbation effects from simpler cell systems to a more complex co-culture system. We have found similar trends in model performance, with more details in Appendix 9.1
Our simpler baselines, in particular the latent additive (LA) model, offer competitive performance at a fraction of training cost. These baseline models also scale well with larger datasets that contain more complex covariates, as shown in the Section on 5.3 with the McFalineFigueroa23 and Jiang24 datasets, where they outperform the more complex VAEs. Thus, it is important to evaluate models across a diverse set of datasets, which we offer in our benchmark.
5.2 Predicting Combinatorial Gene Over-expression Effects
Next we discuss model performance on the combo prediction task using the Norman19 dataset, which contains both single and dual genetic perturbations. The models see all single perturbations and 30% of the dual perturbations during training, with the remaining 70% of duals held out for validation and testing. We summarize the results in Table 3.
The latent additive, decoder-only, and a subset of published models outperform the linear model across all metrics, suggesting that deep learning approaches can capture some non-linear interactions in the Norman19 dataset. However, the linear model performance performance is still relatively strong, suggesting most dual perturbation effects are linear.
Again, removing the adversary component or sparsity constraint leads to better performance, as is the case of CPA (noAdv) and SAMS-VAE (S). Both models perform largely on par with our best baseline latent additive model. Overall, Norman19 offers clean perturbation effects, ideal for sanity checking models during development.
Table 3: Results of the combo prediction experiment. Model performance predicting dual perturbation effects in the Norman19 dataset.
Model Cosine RMSE Cosine RMSE
LogFC mean LogFC rank mean rank
CPA $0.76\pm 4\times 10^{-3}$ $0.046\pm 5\times 10^{-4}$ $0.0072\pm 2\times 10^{-3}$ $0.019\pm 3\times 10^{-3}$
CPA (noAdv) $0.77\pm 1\times 10^{-2}$ $0.046\pm 1\times 10^{-3}$ $\mathbf{0.0057\pm 3\times 10^{-3}}$ $0.017\pm 3\times 10^{-3}$
CPA (scGPT) $0.70\pm 2\times 10^{-2}$ $0.053\pm 1\times 10^{-3}$ $0.025\pm 6\times 10^{-3}$ $0.036\pm 3\times 10^{-3}$
SAMS-VAE $0.45\pm 2\times 10^{-2}$ $0.084\pm 7\times 10^{-4}$ $0.021\pm 5\times 10^{-3}$ $0.026\pm 2\times 10^{-3}$
SAMS-VAE (S) $\mathbf{0.78\pm 6\times 10^{-3}}$ $0.047\pm 2\times 10^{-3}$ $0.019\pm 5\times 10^{-3}$ $0.030±\pm 7\times 10^{-3}$
Biolord $0.41\pm 2\times 10^{-2}$ $0.086\pm 6\times 10^{-4}$ $0.027\pm 1\times 10^{-3}$ $0.028\pm 1\times 10^{-3}$
GEARS $0.44\pm 5\times 10^{-3}$ $0.069\pm 1\times 10^{-3}$ $0.051\pm 1\times 10^{-2}$ $0.055\pm 6\times 10^{-3}$
LA $\mathbf{0.79\pm 1\times 10^{-2}}$ $\mathbf{0.043\pm 4\times 10^{-4}}$ $\mathbf{0.005\pm 2\times 10^{-3}}$ $\mathbf{0.014\pm 1\times 10^{-3}}$
LA (scGPT) $0.77\pm 4\times 10^{-3}$ $\mathbf{0.044\pm 4\times 10^{-4}}$ $0.0085\pm 1\times 10^{-3}$ $\mathbf{0.013\pm 2\times 10^{-3}}$
Decoder $0.73\pm 2\times 10^{-2}$ $\mathbf{0.043\pm 3\times 10^{-4}}$ $0.017\pm 6\times 10^{-3}$ $\mathbf{0.014\pm 4\times 10^{-4}}$
Linear $0.60\pm 2\times 10^{-2}$ $0.057\pm 3\times 10^{-3}$ $0.035\pm 4\times 10^{-3}$ $0.016\pm 8\times 10^{-4}$
5.3 Effect of Data Scaling
Figure 3: Scaling of cosine similarity (left) and its rank (right) with increasing size of data included in the training process ($x$-axis) for several perturbation response models. Points represent results on test data for 5 different seeds, the line represent their average.
Table 4: Results of the Jiang24 experiment which contains 30 different cell states (6 cell lines & 5 cytokine treatments), with 70% of perturbations from 9 cell states held out for validation/testing. Results are reported as mean $\pm$ one standard deviation. Best performance per metric is indicated in bold.
Model Cosine RMSE Cosine RMSE
LogFC mean LogFC rank mean rank
CPA $0.60\pm 7\times 10^{-4}$ $\mathbf{0.015\pm 5\times 10^{-5}}$ $0.40\pm 9\times 10^{-3}$ $0.42\pm 1\times 10^{-2}$
CPA (noAdv) $0.60\pm 2\times 10^{-3}$ $\mathbf{0.015\pm 4\times 10^{-5}}$ $0.39\pm 1\times 10^{-2}$ $0.40\pm 1\times 10^{-2}$
SAMS-VAE $0.59\pm 3\times 10^{-3}$ $0.017\pm 2\times 10^{-4}$ $0.45\pm 9\times 10^{-3}$ $0.45\pm 6\times 10^{-3}$
SAMS-VAE (S) $0.57\pm 5\times 10^{-2}$ $0.016\pm 2\times 10^{-3}$ $0.42\pm 2\times 10^{-2}$ $0.42\pm 3\times 10^{-2}$
LA $0.47\pm 1\times 10^{-3}$ $\mathbf{0.015\pm 7\times 10^{-5}}$ $0.38\pm 6\times 10^{-3}$ $0.38\pm 7\times 10^{-3}$
Decoder $\mathbf{0.64\pm 8\times 10^{-4}}$ $\mathbf{0.015\pm 3\times 10^{-5}}$ $\mathbf{0.32\pm 8\times 10^{-3}}$ $\mathbf{0.32\pm 5\times 10^{-3}}$
Linear $0.17\pm 8\times 10^{-5}$ $0.038\pm 5\times 10^{-5}$ $0.34\pm 2\times 10^{-3}$ $0.43\pm 1\times 10^{-3}$
In this section, we report the data scaling results, assessing whether models can take advantage of additional training data to better generalize perturbation effects across biological states. We use the McFalineFigueroa23 dataset that contains 3 cell lines with 5 chemical perturbations and 525 genetic perturbations. We treat unique cell line and chemical perturbation as a separate biological state, resulting in 15 total states. To test whether adding biological states improves performance, we construct nested subsets of the dataset ($small\subset medium\subset full$), all sharing the same validation and test sets. Each of the subsets contains more biological states (details in Appendix 10.4).
We find all models tend to improve with more training data in both cosine LogFC and its rank metric (see Figure 3), with the exception of the CPA and SAMS-VAE models. As the training data increases, CPA and CPA (noAdv) perform worse on the cosine LogFC metric, while SAMS-VAE and SAMS-VAE (S) perform worse on the rank metric. Our baseline models, especially the simple latent additive model, generally outperform the more complex CPA and SAMS-VAE models, especially on the rank metric and with larger training data sizes. This suggests that the simpler baselines may scale to large datasets better than the more complex published models.
We further assessed how these models perform on large datasets with complex covariates by applying our benchmarking suite to the Jiang24 dataset, a large, 1.6 million cell dataset with complex covariates. The dataset contained 6 cell lines with 5 unique cytokine treatments, which we modeled as 30 distinct biological states and 219 genetic perturbations. However, the set of perturbations applied was different for each cytokine treatment. We used a similar splitting strategy where we held out 70% of the covariates from 9 cell states for validation/testing.
The results are summarized in Table 4 where the decoder-only model performs the best for all metrics considered. Here, none of the published models perform on-par with our simple baselines, particularly on the rank metrics which indicates these more sophisticated models are unable to generate perturbation-specific responses. Again, this aligns with our earlier observation that simpler architectures with fewer modeling assumptions benefit more from additional training data.
6 Discussion
Summary
Our study shows that for predicting average perturbation effects, there is no model, or even class of models, that is clearly better than others. In smaller datasets, methods such as CPA (noAdv) and SAMS-VAE (S) exhibit the best performance. However, these methods do not scale well in larger datasets, where they are outperformed by our latent additive or decoder only models using the matching approach. The variance in performance across datasets highlights the need to develop more versatile and universally robust models and we hope that our perturbation framework will be helpful for developing and benchmarking these future models.
We also find that ablating the adversarial component in CPA and the sparsity-inducing component in SAMS-VAE improve performance over the original models. Since it is highly probable that assumptions made and tested in one dataset may not hold true in another, comprehensive ablation experiments across diverse datasets and metrics as are provided in our framework will be a gateway to developing more versatile and robust model.
Our simple decoder-only models with no gene expression input generally perform well and on Jiang24 dataset, significantly outperforms all other models. This potentially points to new architectures that emphasize the perturbation encoder, assuming the information present in the gene expression input may not be as rich as the perturbation and covariate information. But other possible explanations include inadequate control matching due to heterogeneity in the control cells not captured by covariate labels, as well as noise in either the single-cell RNA readouts or the perturbation labels. Specifically, CRISPR knockdown or knockout can be inefficient, resulting in some cells receiving the correct guide RNA but not having a robust gene knockdown or knockout [33].
The performance on the larger datasets and tasks, that is McFalineFigueroa23 (data scaling) and Jiang24, are unsatisfactory compared to what was observed on the smaller datasets. As an example, on the data scaling task, none of the models got rank metric below $0.4$ (where $0.5$ would be achieved by a random predictor). This could be due to both the intrinsic difficulty of the task, and the amount of measurement or biological noise present in a dataset. For the data scaling task, most models performed better with more data, potentially suggesting perturbation models follow scaling laws [26].
Limitations
Given the diversity of implementations among public models, we aimed to assess the core components of each model. Thus, our benchmarking results should be interpreted as an assessment of how these core components perform rather than a perfectly accurate recreation of the public model implementation (see Appendix 10.5 for details).
Since our benchmarking metrics are defined on a population level, they may not fully capture heterogeneity in the perturbation response among cells. We’ve implemented distributional metrics, such as maximum mean discrepancy (MMD) with RBF kernel and energy distance, as well as common differential expression metrics, such as log p-values, to better capture response heterogeneity and hope to use these metrics in future studies [20, 53].
Benchmarking codebase
We provide three main components: datasets and dataloaders, a model development framework, and an evaluation API with metrics (Appendix 7). Each component can be used together or individually. For example, a user who just wants to benchmark predictions generated by an existing model can use the evaluation API. Whereas a user who wants a develop a new model with our model framework can use the entire codebase. Each component is extensible, making it easy for users to add new datasets, models, and metrics. The codebase can be found on GitHub at https://github.com/altoslabs/perturbench/.
Conclusion
The perturbation response modeling field holds great promise in searching the truly vast space of genetic and chemical perturbations to find disease targets and potential therapeutics. In this work, we bring together state-of-the-art models in a unified framework with thorough evaluation. We benchmarked individual model components through ablation studies, assessed performance on different datasets at different scales, and developed rank metrics that capture key downstream applications and model failure modes. Finally, we hope that our modular codebase will prove valuable in future model development and benchmarking efforts.